OpenStudy (anonymous):

Solve 11.3^(2x – 1) = 15.7

sam (.sam.):

11.3^(2x-1)=15.7 ln(11.3^(2x-1))=ln(15.7) 4.8496x-2.4248=ln(15.7) 4.8496x=5.1785 x=1.0678

OpenStudy (anonymous):

thanksss your are amazinggg